# EPHOU-24-015 KYUSHU-HET-299 Modular symmetry of localized modes

### v1

@@ 1 Introduction
This paper is organized as follows. In section 2, we review a 2D toroidal orbifold compactification. We give a brief review on the construction in subsection 2.1, the modular symmetry and modular forms in subsection 2.2, and localized modes on orbifold fixed points in subsection 2.3. In section 3, we study the modular symmetry of the localized modes. We study the general discussions in subsection 3.1, the specific wave functions of the localized modes in subsection 3.2, and various patterns of localized magnetic fluxes in subsection 3.3. In section 4, we conclude this paper.
@@ 3.1 Modular symmetry
$(z^{\prime},\tau^{\prime})=(-\frac{z}{\tau},-\frac{1}{\tau}).$ (3.1)
@@
$\begin{pmatrix}\psi_{{\rm L},0}(S^{2}(z,\tau))\\ \psi_{{\rm L},1/2}(S^{2}(z,\tau))\\ \psi_{{\rm L},\tau/2}(S^{2}(z,\tau))\\ \psi_{{\rm L},(\tau+1)/2}(S^{2}(z,\tau))\end{pmatrix}=e^{-2\pi i\ell}\begin{pmatrix}s_{11}^{4}&0&0&0\\ 0&s_{23}^{4}&0&0\\ 0&0&s_{23}^{4}&0\\ 0&0&0&s_{44}^{4}\end{pmatrix}\begin{pmatrix}\psi_{{\rm L},0}(z,\tau)\\ \psi_{{\rm L},1/2}(z,\tau)\\ \psi_{{\rm L},\tau/2}(z,\tau)\\ \psi_{{\rm L},(\tau+1)/2}(z,\tau)\end{pmatrix},$ (3.8)
@@
However, both of these relations must correspond to the boundary conditions of wave functions by $2\pi$ rotation, because both Eqs. (3.6) and (3.7) are $2\pi$ rotation. We consider the circle around $z=0$. Then, the phase $e^{2\pi if}$ appears in the above boundary conditions, where $f$ corresponds to the total amount of the magnetic flux inserted inside the circle. When the fixed point at $z=0$ is only contained inside the circle, only $\psi_{{\rm L},0}$ feels a total magnetic flux, $\xi^{F}_{0}=2\ell$. Thus, the boundary conditions lead to the following relations:
$\begin{pmatrix}\psi_{{\rm L},0}(S^{2}(z,\tau))\\ \psi_{{\rm L},1/2}(S^{2}(z,\tau))\\ \psi_{{\rm L},\tau/2}(S^{2}(z,\tau))\\ \psi_{{\rm L},(\tau+1)/2}(S^{2}(z,\tau))\end{pmatrix}=\begin{pmatrix}e^{4\pi i\ell}&0&0&0\\ 0&1&0&0\\ 0&0&1&0\\ 0&0&0&1\end{pmatrix}\begin{pmatrix}\psi_{{\rm L},0}(z,\tau)\\ \psi_{{\rm L},1/2}(z,\tau)\\ \psi_{{\rm L},\tau/2}(z,\tau)\\ \psi_{{\rm L},(\tau+1)/2}(z,\tau)\end{pmatrix},$ (3.10)
@@
Furthermore, under the $T$ transformation, localized modes at $z=0$ and $z=1/2$ do not shift from ones at $z=0$ and $z=1/2$, respectively. Thus, the wave functions $\psi_{{\rm L},0}$ and $\psi_{{\rm L},(\tau+1)/2}$ may change by the same phase, i.e.,
@@ 3.2 Specific wave functions of localized modes
In this subsection, we show specific wave functions of localized modes. We denote the wave function localized around $z=z_{{\rm f.p.}}$ by $\psi_{z_{{\rm f.p.}},\mathbb{Z}_{N}^{1}}(z,\tau)$. It can be approximated around $z=z_{{\rm f.p.}}$ as
@@
Note that $\left(\psi_{z_{{\rm f.p.}},\mathbb{Z}_{N}^{1}}(z,\tau)\right)^{N}$ is $\mathbb{Z}_{N}$ invariant although the $\mathbb{Z}_{N}$ charge of $\psi_{z_{{\rm f.p.}},\mathbb{Z}_{N}^{1}}(z,\tau)$ is $1$. Then, by using $\mathbb{Z}_{N}$ invariant mode, $\psi_{z_{{\rm f.p.}},\mathbb{Z}_{N}^{0}}(z,\tau)$, the $a$-th localized mode can be expressed as [9]
$\psi_{{\rm L}}^{a}(z,\tau)=\left(\psi_{z_{{\rm f.p.}},\mathbb{Z}_{N}^{0}}(z,\tau)/\psi_{z_{{\rm f.p.}},\mathbb{Z}_{N}^{1}}(z,\tau)\right)^{(\ell-a)N}\psi_{{\rm B}}(z,\tau),$ (3.37)
where $\psi_{{\rm B}}(z,\tau)$ denotes the bulk mode obtained by the singular gauge transformation:
@@
On the other hand, there are three candidates of $\psi_{0,\mathbb{Z}_{N}^{0}}(z,\tau)$, i.e., $\psi^{(0,0)}(z,\tau)$, $\psi^{(1/2,0)}(z,\tau)$, and $\psi^{(0,1/2)}(z,\tau)$. Thus, the localized mode at $z_{{\rm f.p.}}=0$ can be written by
$\psi_{{\rm L},0}^{a}(z,\tau)=({\rm Im}\tau)^{-\ell N/4}\left[\psi^{(1/2,1/2)}(z,\tau)/\left|\psi^{(1/2,1/2)}(z,\tau)\right|\right]^{\ell N}\left[\Phi(z,\tau)/\left(\psi^{(1/2,1/2)}(z,\tau)\right)^{N}\right]^{\ell-a},$
$\Phi(z,\tau)=C_{0}^{(0,0)}(\tau)\left(\psi^{(0,0)}(z,\tau)\right)^{N}+C_{0}^{(1/2,0)}(\tau)\left(\psi^{(1/2,0)}(z,\tau)\right)^{N}+C_{0}^{(0,1/2)}(\tau)\left(\psi^{(0,1/2)}(z,\tau)\right)^{N},$ (3.39)
where $C_{0}^{(\alpha_{1},\alpha_{2})}(\tau)$ denote coefficients that depend on not $z$ but $\tau$.
Now, by considering the following modular transformations [35]:
$\begin{array}[]{cc}\psi^{(0,0)}(S(z,\tau))=(-\tau)^{1/2}e^{\pi i/4}\psi^{(0,0)}(z,\tau),&\psi^{(0,0)}(T(z,\tau))=\psi^{(0,1/2)}(z,\tau),\\ \psi^{(1/2,0)}(S(z,\tau))=(-\tau)^{1/2}e^{\pi i/4}\psi^{(0,1/2)}(z,\tau),&\psi^{(1/2,0)}(T(z,\tau))=e^{\pi i/4}\psi^{(1/2,0)}(z,\tau),\\ \psi^{(0,1/2)}(S(z,\tau))=(-\tau)^{1/2}e^{\pi i/4}\psi^{(1/2,0)}(z,\tau),&\psi^{(0,1/2)}(T(z,\tau))=\psi^{(0,0)}(z,\tau),\\ \psi^{(1/2,1/2)}(S(z,\tau))=(-\tau)^{1/2}e^{3\pi i/4}\psi^{(1/2,1/2)}(z,\tau),&\psi^{(1/2,1/2)}(T(z,\tau))=e^{\pi i/4}\psi^{(1/2,1/2)}(z,\tau),\end{array}$ (3.44)
we set the coefficients, $C_{0}^{(\alpha_{1},\alpha_{2})}(\tau)$, such that the localized modes can be linearly transformed under the modular transformation:
@@
Namely, the localized mode at $z_{{\rm f.p.}}=0$ can be expressed as
@@
Note again that since we chose the coefficients as Eq. (3.48), all of the three terms of $\Phi(z,\tau)$ in Eq. (3.39) have the same behavior under the modular transformation. Also, the form of the wave function (3.49) has the following permutation symmetry:
@@
Similarly, we can derive the localized modes at the other fixed points, $z=z_{\rm f.p.}$, by replacing $z$ with $Z_{z_{\rm f.p.}}=z-z_{\rm f.p.}$ and considering the following relations:
$\begin{array}[]{l}\psi^{(0,0)}(Z_{1/2},\tau)=e^{i\chi_{1}}\psi^{(0,1/2)}(z,\tau)\\ \psi^{(1/2,0)}(Z_{1/2},\tau)=e^{i\chi_{1}}e^{-\pi i/2}\psi^{(1/2,1/2)}(z,\tau)\\ \psi^{(0,1/2)}(Z_{1/2},\tau)=e^{i\chi_{1}}\psi^{(0,0)}(z,\tau)\\ \psi^{(1/2,1/2)}(Z_{1/2},\tau)=e^{i\chi_{1}}e^{-\pi i/2}\psi^{(1/2,0)}(z,\tau)\end{array},$ (3.57)
$\begin{array}[]{l}\psi^{(0,0)}(Z_{\tau/2},\tau)=e^{i\chi_{2}}\psi^{(1/2,0)}(z,\tau)\\ \psi^{(1/2,0)}(Z_{\tau/2},\tau)=e^{i\chi_{2}}\psi^{(0,0)}(z,\tau)\\ \psi^{(0,1/2)}(Z_{\tau/2},\tau)=e^{i\chi_{2}}e^{-\pi i}\psi^{(1/2,1/2)}(z,\tau)\\ \psi^{(1/2,1/2)}(Z_{\tau/2},\tau)=e^{i\chi_{2}}\psi^{(0,1/2)}(z,\tau)\end{array},$ (3.62)
$\begin{array}[]{l}\psi^{(0,0)}(Z_{(\tau+1)/2},\tau)=e^{i\chi_{1+2}}e^{-3\pi i/4}\psi^{(1/2,1/2)}(z,\tau)\\ \psi^{(1/2,0)}(Z_{(\tau+1)/2},\tau)=e^{i\chi_{1+2}}e^{-\pi i/4}\psi^{(0,1/2)}(z,\tau)\\ \psi^{(0,1/2)}(Z_{(\tau+1)/2},\tau)=e^{i\chi_{1+2}}e^{\pi i/4}\psi^{(1/2,0)}(z,\tau)\\ \psi^{(1/2,1/2)}(Z_{(\tau+1)/2},\tau)=e^{i\chi_{1+2}}e^{-\pi i/4}\psi^{(0,0)}(z,\tau)\end{array}.$ (3.67)
Hence, the localized modes at $z=1/2$, $\tau/2$, and $(\tau+2)/2$ can be expressed as
@@
respectively. Note that coefficients $C_{z_{\rm f.p.}}^{(\alpha_{1},\alpha_{2})}=\psi^{(\alpha^{\prime}_{1},\alpha^{\prime}_{2})}(0,\tau)\psi^{(\alpha^{\prime\prime}_{1},\alpha^{\prime\prime}_{2})}(0,\tau)$ are also shifted and then we set $z=0$. They transform under $S$ and $T$ transformations as
@@
Note again that the above choice of coefficients $C_{z_{\rm f.p.}}^{(\alpha_{1},\alpha_{2})}$ are important to lead to the modular transformation behavior, and there is permutation symmetries similar to Eq. (3.52) .
@@
which correspond to Eq. (3.18). Namely, the localized modes have the following modular symmetry:
@@ 3.3 Different localized magnetic fluxes
So far, we have studied the case that the same size of magnetic fluxes $\xi^{F}_{z_{\rm f.p.}}$ are localized on all of the fixed points. Here, we discuss the cases that magnetic fluxes localized at fixed points are different from each other. Note that the localized mode at $z=0$ is a singlet. Thus, when we set only the magnetic flux localized at $z=0$ which is different from the others and the others are the same, the flavor group is the same, i.e., $S_{3}$ or $S^{\prime}_{4}$. However, only the representation of localized mode at $z=0$ changes. For example, when $\ell\in 2(2\mathbb{Z}+1)$ and $4\mathbb{Z}$ for $z=0$, the localized mode at $z=0$ is the non-trivial and trivial singlets of $S_{3}$, respectively, and it is $\mathbb{Z}_{2}$-odd and -even mode where $\mathbb{Z}_{2}$ is a subgroup of $S_{3}$. On the other hand, when $\ell\in 2\mathbb{Z}+1$, the localized mode at $z=0$ has a $\mathbb{Z}_{4}$ charge, where $\mathbb{Z}_{4}$ is a subgroup of $S^{\prime}_{4}$.
@@
We have discussed flavor symmetries by several patterns of localized magnetic fluxes. Certain patterns may be constrained by some reasons. In Refs. [5, 8], it was found that non-vanishing localized fluxes modify the profiles of bulk zero mods and massive Kaluza-Klein modes. Modified profiles of these modes induce radiative corrections on localized fluxes. Then, such radiative corrections modify profiles of bulk modes further. Thus, this system may be unstable. One of stable configurations is the pattern that the sizes of localized magnetic fluxes at all of the fixed points are the same [8]. Therefore, such a pattern of localized fluxes is favorable from the viewpoint of stability.

### v2

@@ 1 Introduction
This paper is organized as follows. In section 2, we review a 2D toroidal orbifold compactification. We give a brief review on the construction in subsection 2.1, the modular symmetry and modular forms in subsection 2.2, and localized modes on orbifold fixed points in subsection 2.3. In section 3, we study the modular symmetry of the localized modes. We study the general discussions in subsection 3.1, the specific wave functions of the localized modes in subsection 3.2, and various patterns of localized magnetic fluxes in subsection 3.3. In section 4, we conclude this paper. In Appendix A, we summarize useful relations of wave functions, discussed in subsection 3.1.
@@ 3.1 Modular symmetry
$(z^{\prime},\tau^{\prime})=\left(-\frac{z}{\tau},-\frac{1}{\tau}\right).$ (3.1)
@@
$\begin{pmatrix}\psi_{{\rm L},0}(S^{4}(z,\tau))\\ \psi_{{\rm L},1/2}(S^{4}(z,\tau))\\ \psi_{{\rm L},\tau/2}(S^{4}(z,\tau))\\ \psi_{{\rm L},(\tau+1)/2}(S^{4}(z,\tau))\end{pmatrix}=e^{-2\pi i\ell}\begin{pmatrix}s_{11}^{4}&0&0&0\\ 0&s_{23}^{4}&0&0\\ 0&0&s_{23}^{4}&0\\ 0&0&0&s_{44}^{4}\end{pmatrix}\begin{pmatrix}\psi_{{\rm L},0}(z,\tau)\\ \psi_{{\rm L},1/2}(z,\tau)\\ \psi_{{\rm L},\tau/2}(z,\tau)\\ \psi_{{\rm L},(\tau+1)/2}(z,\tau)\end{pmatrix},$ (3.8)
@@
However, both of these relations must correspond to the boundary conditions of wave functions by $2\pi$ rotation, because both Eqs. (3.6) and (3.7) are $2\pi$ rotation. We consider the circle around $z=0$. Then, the phase $e^{2\pi if}$ appears in the above boundary conditions, where $f$ corresponds to the total amount of the magnetic flux inserted inside the circle. [^2] Precisely, $f$ denotes the total amount of the effective magnetic flux including the effect of the localized curvature at the orbifold fixed point. In the following analysis, we consider the effective magnetic fluxes. When the fixed point at $z=0$ is only contained inside the circle, only $\psi_{{\rm L},0}$ feels a total magnetic flux, $\xi^{F}_{0}=2\ell$. Thus, the boundary conditions lead to the following relations:
$\begin{pmatrix}\psi_{{\rm L},0}(S^{4}(z,\tau))\\ \psi_{{\rm L},1/2}(S^{4}(z,\tau))\\ \psi_{{\rm L},\tau/2}(S^{4}(z,\tau))\\ \psi_{{\rm L},(\tau+1)/2}(S^{4}(z,\tau))\end{pmatrix}=\begin{pmatrix}e^{4\pi i\ell}&0&0&0\\ 0&1&0&0\\ 0&0&1&0\\ 0&0&0&1\end{pmatrix}\begin{pmatrix}\psi_{{\rm L},0}(z,\tau)\\ \psi_{{\rm L},1/2}(z,\tau)\\ \psi_{{\rm L},\tau/2}(z,\tau)\\ \psi_{{\rm L},(\tau+1)/2}(z,\tau)\end{pmatrix},$ (3.10)
@@
Furthermore, under the $T$ transformation, localized modes at $z=0$ and $z=1/2$ do not shift from ones at $z=0$ and $z=1/2$, respectively. Thus, the wave functions $\psi_{{\rm L},0}$ and $\psi_{{\rm L},1/2}$ may change by the same phase, i.e.,
@@ 3.2 Specific wave functions of localized modes
In this subsection, we show specific wave functions of localized modes. As in Ref. [9], they can be obtained by modifying the wave function of the bulk mode after the singular gauge transformation. Therefore, we first review the singular gauge transformation around an orbifold fixed point, $z=z_{{\rm f.p.}}$, and the wave function of the bulk mode after the singular gauge transformation in 3.2.1. Then, we show specific wave function of the localized modes, and study their modular flavor symmetry in 3.2.2.
3.2.1 Wave function of the bulk mode after the singular gauge transformation
To define the singular gauge transformation around an orbifold fixed point, $z=z_{{\rm f.p.}}$, we introduce the wave function, $\psi_{z_{{\rm f.p.}},\mathbb{Z}_{N}^{1}}(z,\tau)$, such that it can be approximated around $z=z_{{\rm f.p.}}$ as
@@ 3.2.1 Wave function of the bulk mode after the singular gauge transformation
Then, we can express the bulk mode after the singular gauge transformation around the fixed point, $z=z_{{\rm f.p.}}$, as
@@
3.2.2 Wave functions of localized modes and their modular symmetry
Now, let us show wave functions of localized modes by modifying the wave function of the bulk mode in Eq. (3.37), and study their modular flavor symmetry. Here, we note that $\left(\psi_{z_{{\rm f.p.}},\mathbb{Z}_{N}^{1}}(z,\tau)\right)^{N}$ in Eq. (3.37) is $\mathbb{Z}_{N}$ invariant although the $\mathbb{Z}_{N}$ charge of $\psi_{z_{{\rm f.p.}},\mathbb{Z}_{N}^{1}}(z,\tau)$ is $1$. Thus, by introducing the $\mathbb{Z}_{N}$ invariant mode, $\psi_{z_{{\rm f.p.}},\mathbb{Z}_{N}^{0}}(z,\tau)$, and replacing $\left(\psi_{z_{{\rm f.p.}},\mathbb{Z}_{N}^{1}}(z,\tau)\right)^{N}$ with $\left(\psi_{z_{{\rm f.p.}},\mathbb{Z}_{N}^{0}}(z,\tau)\right)^{N}$, the $a$-th localized mode can be expressed as [9]
$\psi_{{\rm L}}^{a}(z,\tau)=\left(\psi_{z_{{\rm f.p.}},\mathbb{Z}_{N}^{0}}(z,\tau)/\psi_{z_{{\rm f.p.}},\mathbb{Z}_{N}^{1}}(z,\tau)\right)^{(\ell-a)N}\psi_{{\rm B}}(z,\tau).$ (3.38)
From now, we discuss the specific wave function of $\psi_{z_{{\rm f.p.}},\mathbb{Z}_{N}^{0}}(z,\tau)$. First, let us consider $z_{{\rm f.p.}}=0$ case. In this case, there are three candidates of $\psi_{0,\mathbb{Z}_{N}^{0}}(z,\tau)$, i.e., $\psi^{(0,0)}(z,\tau)$, $\psi^{(1/2,0)}(z,\tau)$, and $\psi^{(0,1/2)}(z,\tau)$. Thus, the localized mode at $z_{{\rm f.p.}}=0$ can be written by
$\psi_{{\rm L},0}^{a}(z,\tau)=({\rm Im}\tau)^{-\ell N/4}\left[\psi^{(1/2,1/2)}(z,\tau)/\left|\psi^{(1/2,1/2)}(z,\tau)\right|\right]^{\ell N}\left[\Phi_{0}(z,\tau)/\left(\psi^{(1/2,1/2)}(z,\tau)\right)^{N}\right]^{\ell-a},$
$\Phi_{0}(z,\tau)=C_{0}^{(0,0)}(\tau)\left(\psi^{(0,0)}(z,\tau)\right)^{N}+C_{0}^{(1/2,0)}(\tau)\left(\psi^{(1/2,0)}(z,\tau)\right)^{N}+C_{0}^{(0,1/2)}(\tau)\left(\psi^{(0,1/2)}(z,\tau)\right)^{N},$ (3.39)
where $C_{0}^{(\alpha_{1},\alpha_{2})}(\tau)$ denote coefficients that depend on not $z$ but $\tau$. In particular, in order for the localized modes to be linearly transformed under the modular transformation, we set the coefficients, $C_{0}^{(\alpha_{1},\alpha_{2})}(\tau)$, as
@@ 3.2.2 Wave functions of localized modes and their modular symmetry
Here, we consider the modular transformation of $\psi^{(\alpha_{1},\alpha_{2})}(z,\tau)$, summarized in Eq. (A.5). Namely, the localized mode at $z_{{\rm f.p.}}=0$ can be expressed as
@@
Note again that since we chose the coefficients as Eq. (3.43), all of the three terms of $\Phi_{0}(z,\tau)$ in Eq. (3.39) have the same behavior under the modular transformation. Also, the form of the wave function (3.44) has the following permutation symmetry:
@@
Similarly, by replacing $z$ with $Z_{z_{\rm f.p.}}=z-z_{\rm f.p.}$ and considering the relations in Eqs. (A.10)-(A.20), the localized modes at the other fixed points, $z=1/2$, $\tau/2$, and $(\tau+1)/2$, can be expressed as
@@
respectively. Note that coefficients $C_{z_{\rm f.p.}}^{(\alpha_{1},\alpha_{2})}=\psi^{(\alpha^{\prime}_{1},\alpha^{\prime}_{2})}(0,\tau)\psi^{(\alpha^{\prime\prime}_{1},\alpha^{\prime\prime}_{2})}(0,\tau)$ are also shifted by $z_{\rm f.p.}$ and then we set $z=0$. They transform under the $S$ and $T$ transformations as
@@
Note again that the above choice of coefficients $C_{z_{\rm f.p.}}^{(\alpha_{1},\alpha_{2})}$ is important to lead to the modular transformation behavior, and there are permutation symmetries similar to Eq. (3.47) .
@@
which correspond to Eq. (3.18). Namely, the localized modes have the following modular flavor symmetry:
@@ 3.3 Different localized magnetic fluxes
So far, we have studied the case that the same sizes of magnetic fluxes $\xi^{F}_{z_{\rm f.p.}}$ are localized on all of the fixed points. Here, we discuss the cases that magnetic fluxes localized at fixed points are different from each other. Note that the localized mode at $z=0$ is a singlet. Thus, when we set only the magnetic flux localized at $z=0$ which is different from the others and the others are the same, the flavor group is the same, i.e., $S_{3}$ or $S^{\prime}_{4}$. However, only the representation of localized mode at $z=0$ changes. For example, when $\ell\in 2(2\mathbb{Z}+1)$ and $4\mathbb{Z}$ for $z=0$, the localized mode at $z=0$ is the non-trivial and trivial singlets of $S_{3}$, respectively, and it is $\mathbb{Z}_{2}$-odd and -even mode where $\mathbb{Z}_{2}$ is a subgroup of $S_{3}$. On the other hand, when $\ell\in 2\mathbb{Z}+1$, the localized mode at $z=0$ has a $\mathbb{Z}_{4}$ charge, where $\mathbb{Z}_{4}$ is a subgroup of $S^{\prime}_{4}$.
@@
We have discussed flavor symmetries by several patterns of localized magnetic fluxes. Certain patterns may be constrained by some reasons. The degree of freedom of localized fluxes is equivalent to the Fayet-Iliopoulos (FI) terms localized at fixed points [5]. In Refs. [5, 8], it was found that non-vanishing FI terms modify the profiles of bulk zero modes and massive Kaluza-Klein modes. Modified profiles of these modes induce radiative corrections on FI terms. Then, such radiative corrections modify profiles of bulk modes further. Thus, this system may be unstable. One of stable configurations is the pattern that the sizes of FI terms at all of the fixed points are the same [8]. Therefore, such a pattern of FI terms localized fluxes is favorable from the viewpoint of stability.
@@ Acknowledgments
Appendix A Summary of useful relations of wave functions
First, we summarize the modular transformation of wave functions, $\psi^{(\alpha_{1},\alpha_{2})}(z,\tau)$ [35]. Especially, they transform under $S$ and $T$ transformations as
$\begin{array}[]{cc}\psi^{(0,0)}(S(z,\tau))=(-\tau)^{1/2}e^{\pi i/4}\psi^{(0,0)}(z,\tau),&\psi^{(0,0)}(T(z,\tau))=\psi^{(0,1/2)}(z,\tau),\\ \psi^{(1/2,0)}(S(z,\tau))=(-\tau)^{1/2}e^{\pi i/4}\psi^{(0,1/2)}(z,\tau),&\psi^{(1/2,0)}(T(z,\tau))=e^{\pi i/4}\psi^{(1/2,0)}(z,\tau),\\ \psi^{(0,1/2)}(S(z,\tau))=(-\tau)^{1/2}e^{\pi i/4}\psi^{(1/2,0)}(z,\tau),&\psi^{(0,1/2)}(T(z,\tau))=\psi^{(0,0)}(z,\tau),\\ \psi^{(1/2,1/2)}(S(z,\tau))=(-\tau)^{1/2}e^{3\pi i/4}\psi^{(1/2,1/2)}(z,\tau),&\psi^{(1/2,1/2)}(T(z,\tau))=e^{\pi i/4}\psi^{(1/2,1/2)}(z,\tau),\end{array}$ (A.5)
Next, we summarize the relations of wave functions, $\psi^{(\alpha_{1},\alpha_{2})}$, at $z$ and $Z_{z_{\rm f.p.}}=z-z_{\rm f.p.}$ as follows,
$\begin{array}[]{l}\psi^{(0,0)}(Z_{1/2},\tau)=e^{i\chi_{1}}\psi^{(0,1/2)}(z,\tau)\\ \psi^{(1/2,0)}(Z_{1/2},\tau)=e^{i\chi_{1}}e^{-\pi i/2}\psi^{(1/2,1/2)}(z,\tau)\\ \psi^{(0,1/2)}(Z_{1/2},\tau)=e^{i\chi_{1}}\psi^{(0,0)}(z,\tau)\\ \psi^{(1/2,1/2)}(Z_{1/2},\tau)=e^{i\chi_{1}}e^{-\pi i/2}\psi^{(1/2,0)}(z,\tau)\end{array},$ (A.10)
$\begin{array}[]{l}\psi^{(0,0)}(Z_{\tau/2},\tau)=e^{i\chi_{2}}\psi^{(1/2,0)}(z,\tau)\\ \psi^{(1/2,0)}(Z_{\tau/2},\tau)=e^{i\chi_{2}}\psi^{(0,0)}(z,\tau)\\ \psi^{(0,1/2)}(Z_{\tau/2},\tau)=e^{i\chi_{2}}e^{-\pi i}\psi^{(1/2,1/2)}(z,\tau)\\ \psi^{(1/2,1/2)}(Z_{\tau/2},\tau)=e^{i\chi_{2}}\psi^{(0,1/2)}(z,\tau)\end{array},$ (A.15)
$\begin{array}[]{l}\psi^{(0,0)}(Z_{(\tau+1)/2},\tau)=e^{i\chi_{1+2}}e^{-3\pi i/4}\psi^{(1/2,1/2)}(z,\tau)\\ \psi^{(1/2,0)}(Z_{(\tau+1)/2},\tau)=e^{i\chi_{1+2}}e^{-\pi i/4}\psi^{(0,1/2)}(z,\tau)\\ \psi^{(0,1/2)}(Z_{(\tau+1)/2},\tau)=e^{i\chi_{1+2}}e^{\pi i/4}\psi^{(1/2,0)}(z,\tau)\\ \psi^{(1/2,1/2)}(Z_{(\tau+1)/2},\tau)=e^{i\chi_{1+2}}e^{-\pi i/4}\psi^{(0,0)}(z,\tau)\end{array}.$ (A.20)